# The Quantum Complexity of String Breaking in the Schwinger Model

### v1

@@ Appendix C Charge Rearrangement During String Breaking
Figure 11: The charge sector weights $p_{Q}$ as a function of external charge separation $d$ for the two different orientations of the external field, $E_{\text{ext}}=$- and $E_{\text{ext}}=+$. The system parameters $N=220,a=1,m_{\text{lat}}=0.045,g=0.09$ are used as in Table 1.
@@ Appendix E Supplemental Quantum Complexity Results
Figure 14: The dependence of the RoM on system size and lattice spacing. Top: The vacuum-subtracted RoM of adjacent sites, $\Delta R$ for a selection of system sizes and lattice spacings $(N,a)=(220,1)$ (left), $(440,1/2)$ (center), and $(880,1/4)$ (right). The RoM is plotted in units of $a$ for half of the lattice $n\geq N/2$ as a function of external charge separation $d$. The system parameters from Table 1 are used. Middle: the associated charge densities for half of the lattice $n\leq N/2$. Bottom: the RoM of disjoint regions, $\Delta R_{AB}$, with region AA at the center of the lattice and region $B$ a distance $r$ away. The vacuum value ($d=0$) and the value as $r\to\infty$ are subtracted. The RoM is plotted in units of $a$ for the same system parameters.
@@
Figure 15 details the behavior of the quantum complexity as $m/g$ is varied. We consider several values of $m$ holding $g$ fixed as in the main text, and expand the lattice to account for the increased correlation length with smaller mass. The top and middle rows, as in Fig. 6, display the RoM between adjacent sites, $\Delta R$, and the charge density $q_{n}$ as a function of position $n$ and separation $d$. The transition from strings to mesons becomes increasingly sharper with larger $m/g$, which is seen in both $\Delta R$ and $q_{n}$. The string also appears to break at a larger separation $d$ for larger $m/g$. This is an expected result of fermions being more “expensive” to excite from the vacuum to screen the external charges. The bottom row, as in Fig. 14, shows the RoM between disjoint $\Delta R_{AB}$ subsystems on the lattice as a function of the distance between the regions $r$. Region $A$ is at the center of the lattice, and region $B$ is distance $r$ away. The vacuum-subtracted $\Delta R_{AB}$ only shows structure at small $r$ for the massless case, consistent with the system only having short-range quantum correlations before string breaking. At intermediate $m/g$, the correlations before string breaking are longest-range, and then die off again at $m/g=1$. The propagation of the bound hadrons is seen in all parameters besides the massless case.
@@ Appendix F Computational Methods
The NL RoM is calculated as the minimum of 100 optimizations as in Eq. (10), each with randomly chosen initial conditions.

### v2

@@ Appendix C Charge Rearrangement During String Breaking
Figure 9: The charge sector weights $p_{Q}$ as a function of external charge separation $d$ for the two different orientations of the external field, $E_{\text{ext}}=$- and $E_{\text{ext}}=+$. The system parameters $N=220,a=1,m_{\text{lat}}=0.045,g=0.09$ are used.
@@ Appendix E Supplemental Quantum Complexity Results
Figure 12: The dependence of the RoM on system size and lattice spacing. Top: The vacuum-subtracted RoM of adjacent sites, $\Delta R$ for a selection of system sizes and lattice spacings $(N,a)=(220,1)$ (left), $(440,1/2)$ (center), and $(880,1/4)$ (right). The RoM is plotted in units of $a$ for half of the lattice $n\geq N/2$ as a function of external charge separation $d$. The system parameters from the main text are used. Middle: the associated charge densities for half of the lattice $n\leq N/2$. Bottom: the RoM of disjoint regions, $\Delta R_{AB}$, with region AA at the center of the lattice and region $B$ a distance $r$ away. The vacuum value ($d=0$) and the value as $r\to\infty$ are subtracted. The RoM is plotted in units of $a$ for the same system parameters.
@@
Figure 13 details the behavior of the quantum complexity as $m/g$ is varied. We consider several values of $m$ holding $g$ fixed as in the main text, and expand the lattice to account for the increased correlation length with smaller mass. The top and middle rows display the RoM between adjacent sites, $\Delta R$, and the charge density $q_{n}$ as a function of position $n$ and separation $d$. The transition from strings to mesons becomes increasingly sharper with larger $m/g$, which is seen in both $\Delta R$ and $q_{n}$. The string also appears to break at a larger separation $d$ for larger $m/g$. This is an expected result of fermions being more “expensive” to excite from the vacuum to screen the external charges. The bottom row, as in Fig. 12, shows the RoM between disjoint $\Delta R_{AB}$ subsystems on the lattice as a function of the distance between the regions $r$. Region $A$ is at the center of the lattice, and region $B$ is distance $r$ away. The vacuum-subtracted $\Delta R_{AB}$ only shows structure at small $r$ for the massless case, consistent with the system only having short-range quantum correlations before string breaking. At intermediate $m/g$, the correlations before string breaking are longest-range, and then die off again at $m/g=1$. The propagation of the bound hadrons is seen in all parameters besides the massless case.
@@ Appendix F Computational Methods
The NL RoM is calculated as the minimum of 1000 optimizations as in Eq. (15), each with randomly chosen initial conditions.